OpenStudy (anonymous):

Distributive property. help please 2(7v + 5) + 4v = -2 - 2(6 - 5v)

OpenStudy (cwrw238):

distributive property: a(b + c) = ab + ac

OpenStudy (cwrw238):

so applying it to the first part of your equation 2(7v + 5) = 2*7v + 2*5 = 14v + 10

OpenStudy (cwrw238):

now apply the same process to -2(6 - 5v):- -2(6 - 5v) = -2*6 -2*-5v can you continue?

OpenStudy (anonymous):

Oohh Okey

OpenStudy (anonymous):

No

OpenStudy (cwrw238):

you need more help?

OpenStudy (anonymous):

Nope ive got it :) Thanks

OpenStudy (cwrw238):

good